OpenStudy (x3_drummerchick):

help! will give medals- how do i start to solve this trig equation? (tan θ − 2)(9 sin^2θ − 1) = 0

OpenStudy (baru):

you can try \(tan \theta -2 =0\) and \(9sin^2\theta -1 \)

OpenStudy (x3_drummerchick):

\[(\tan \theta-2)(9\sin ^{2}\theta-1)\]

OpenStudy (x3_drummerchick):

i thought about splitting them up, but then when i solve them, so i put both functions in my calculator? i got up to this point:

OpenStudy (x3_drummerchick):

tanx=2 and sinx=1/3

OpenStudy (baru):

sorry, i meant \(9sin^2\theta -1\)=0

OpenStudy (baru):

yes exactly

OpenStudy (x3_drummerchick):

so do i plug them into my calculator now, or how d i go about using that info?

OpenStudy (baru):

yeah, your calculator sould have a \(sin^{-1}\) function

OpenStudy (x3_drummerchick):

yes, my text is saying there are 5 different answers though

OpenStudy (baru):

sinx can be +1/3 or -1/3

OpenStudy (x3_drummerchick):

thats where i got confused

OpenStudy (x3_drummerchick):

ohh okay

OpenStudy (baru):

well i can see three solutions lying within 0-360 degrees

OpenStudy (x3_drummerchick):

it has something to do with the solution +pi k, then solution + 2pi K

OpenStudy (anonymous):

\[\sin ^2\theta=\frac{ 1 }{ 9 },2 \sin ^2\theta=\frac{ 2 }{ 9 },1-\cos 2\theta=\frac{ 2 }{ 9 }\] \[\cos 2\theta=1-\frac{ 2 }{ 9 }=\frac{ 7 }{ 9 }\]

OpenStudy (x3_drummerchick):

are those double angle formulas?

OpenStudy (anonymous):

yes

OpenStudy (baru):

sorry didnt get what you meant by +pi k .... between 0-360 there are ...many actually

OpenStudy (x3_drummerchick):

i dont understand why i would need double angle formulas for it

OpenStudy (x3_drummerchick):

here is the example problem:

OpenStudy (anonymous):

\[\cos \alpha=\cos \left( 2n \pi \pm \alpha \right) \]

OpenStudy (baru):

yea, i see it, the first solution is for tan, the next four are for sin, lying in each of the four quadrants

OpenStudy (x3_drummerchick):

ohh okay, but how do i do the sin of 1/3?

OpenStudy (x3_drummerchick):

or, rather, how would i go about getting those values?

OpenStudy (baru):

use wolframalpha

OpenStudy (x3_drummerchick):

huh?

OpenStudy (anonymous):

for sin theta= 1/3 1st and 2nd quadrant, theta and pi-theta for sin theta =-1/3' pi+theta and 2 pi- theta

OpenStudy (baru):

http://www.wolframalpha.com/

OpenStudy (x3_drummerchick):

were not allowed to use calculators, so i dont think i should use it

OpenStudy (baru):

ohh

OpenStudy (x3_drummerchick):

:( he makes us do everything manually

OpenStudy (baru):

wow, this question will take quite a while then

OpenStudy (x3_drummerchick):

actually, the answers are decimals, we must have to use a calculator. that wouldnt make sense. ugh. i hate trig.

OpenStudy (baru):

and i just realized i have no clue how to find inverse manually xD

OpenStudy (x3_drummerchick):

i dont even know. this crap is tough and he gave us 2 examples in class- which both were really basic. he throws stupid hard ones on the homework that have like, 5 solutions and shenanigans..

OpenStudy (x3_drummerchick):

im just gonna email him

OpenStudy (x3_drummerchick):

thank you for your help though!

OpenStudy (baru):

sure :)